OpenStudy (anonymous):

Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 4x + y2 − 6y = −4

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (anonymous):

do you know the circle equation?

OpenStudy (here_to_help15):

They keep beating me grr

OpenStudy (anonymous):

yeah one sec i have it in my notes

OpenStudy (anonymous):

(x-h)^2 + (y-k)^2 = r^2 where (h,k) is the center and r is the radius

OpenStudy (here_to_help15):

Here are your answers now for the work :) Center:: (-2,3) Radius:: (3)

OpenStudy (here_to_help15):

@gaos finish what you started :)

OpenStudy (here_to_help15):

OK so wanna plug them in with me?

OpenStudy (anonymous):

(x^2+4x+4)-4+(y^2-6y+9)-9=-4 (x+2)^2+(y-3)^2=9 (x+2)^2+(y-3)^2=3^2 Circles follow the form: (x-a)+(y-b)=r^2 Therefore the center of the circle is at (-2,3) and the radius is 3.

OpenStudy (anonymous):

got it i think!!

OpenStudy (here_to_help15):

And you continue to suprise me :)

OpenStudy (here_to_help15):

Surprise *

OpenStudy (anonymous):

haha thank you :p

OpenStudy (here_to_help15):

:)